OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (anonymous):

okay the slope intercept form is y=mx+b so to find the slope(m)=y2-y1/ x2-x1 =(-3-6)/(0-3)=-9/-3=3 to find b, choose one of the coordinates and plug into y=mx+b i choose(3,6), we know that (x,y) so plug in 6=3(3)+b, solve for b: b=6-9=-3 so you have m=3 and b=-3 plug into y=mx+b y=3x-3 :)